# High-order Computation of Floquet Multipliers and Subspaces using Multistep Methods

### v1

@@ 6 Algorithm and Complexity Analysis
We present pTOAR, an iterative method for computing the dominant $k$ eigenpairs of the periodic-PEP eq.˜18. It combines the periodic Arnoldi process [12] with TOAR-style basis compression [24] to reduce memory and computational cost, making it suitable for large-scale problems. The remainder of this section details the algorithmic steps and its underlying principles.
The periodic Krylov-Schur (pKS) algorithm [12] is well-suited for the linearized periodic-PEP eq.˜22. It avoids explicitly forming the full product $F_{i+p,i}=L^{(p)}\cdots\allowbreak L^{(1)}$, while maintaining accuracy even for eigenvalues spanning several orders of magnitude. The pKS algorithm constructs a periodic Arnoldi process:
$L^{(i)}V^{(i-1)}=V^{(i)}T^{(i)},$ (45)
@@
We explicitly substitute the block form of $L^{(i)}$ into eq.˜45, and partition the $V^{(i)}$ following the block structure of $L^{(i)}$, as shown in eq.˜46:
$\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}\begin{bmatrix}V_{1}^{(i-1)}\\ \vdots\\ V_{d-1}^{(i-1)}\\ V_{d}^{(i-1)}\end{bmatrix}=\begin{bmatrix}V_{1}^{(i)}\\ \vdots\\ V_{d-1}^{(i)}\\ V_{d}^{(i)}\end{bmatrix}T^{(i)}.$ (46)
The key observation of pTOAR from eq.˜46 is an information overlap relationship across adjacent basis matrices:
@@
where $v^{(i-1)}$ is the last column of $V^{(i-1)}$, and $v_{j}^{(i-1)}$ is its $j$-th block of $v^{(i-1)}$, also the last column of $V_{j}^{(i-1)}$ in eq.˜46.
@@
At last, we evaluate memory requirement for storing the input data. For each $L^{(i)}$, it suffices to store the sampled matrices $G^{(i)}$ instead of the coefficient matrix $A_{j}^{(i)}$ in eq.˜14 and the full $nd\times nd$ matrix $L^{(i)}$. This leads to a memory cost of $n^{\gamma}$ per slice, where $\gamma=1$ if the matrices $A_{j}^{(i)}$ are sparse, and $\gamma=2$ if dense. Combining this with the basis memory and the Ritz extraction memory, the overall memory cost is thus $pn^{\gamma}+pnkd$ under assumption $k\ll n$.
@@
1:
2:Time samples $\{t_{i}\}_{i=0}^{p}$ with $0=t_{0}<t_{1}<\cdots<t_{p}=T$, where $T$ is the period;
3:Discrete evaluations $\{G(t_{i})\}_{i=0}^{p}$ of the system matrix;
4:Target number of eigenpairs $k$.
5:
6:Dominant $k$ Floquet multipliers and their corresponding eigenvectors.
7:Discretization: At each time $t_{i}$, define $\left\{A_{j}^{(i)}\right\}_{j=0}^{d}$ in eq.˜14 implicitly via matrix-vector products routines, and invoke them to define matrix-vector products of $L^{(i)}$, as ineq.˜21.
8:Eigenpair Computation: Apply the pTOAR to compute the dominant $k$ eigenpairs of the linearized periodic EVP eq.˜22.
@@ 7 Numerical Results
This section presents three numerical experiments designed to validate our proposed method. The first is an artificial example constructed from known Floquet eigenpairs to verify convergence theory and the distribution of Spurious eigenvalues. The second is a benchmark problem drawn from [25, 6], representative of scenario 1. We compare our pTOAR method with existing packages MATCONT [10] and AUTO-07p [9]. The third experiment targets an industrial circuit application, representative of scenario 2, where fixed sampling makes collocation infeasible and highlights the advantage of our multistep-based method. All numerical experiments are conducted using MATLAB 2024b on a system equipped with an Intel Core i5-1135G7 CPU at 2.4GHz and 16GB of RAM.
@@
$e_{\text{val}}=\frac{\left|\lambda-e^{\mu T}\right|}{\left|e^{\mu T}\right|},\quad e_{\text{vec}}=\frac{\|x_{p}-x(T)\|_{2}}{\|x(T)\|_{2}},$ (47)
@@
$e_{\text{subspace}}=\Theta(U,V),$ (48)
@@
7.1 Convergence Behavior of Floquet multipliers and Spurious Eiegnvalues
Our first experiment is based on a two-dimensional LPTV system given by eq.˜49,
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\alpha\cos^{2}t+(\beta-2)\sin^{2}t&1+(\beta-\alpha-2)\sin t\cos t\\ -1+(\beta-\alpha-2)\sin t\cos t&\alpha\sin^{2}t+(\beta-2)\cos^{2}t\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$. (49)
@@ 7.1 Convergence Behavior of Floquet multipliers and Spurious Eiegnvalues
$U(t)=\begin{bmatrix}\cos t&\sin t\\ -\sin t&\cos t\end{bmatrix}$. (50)
We discretize one period $0,2\pi]$ with variable stepsize, and apply Gear2, Gear3 methods to obtain the periodic-EVP eq.˜22. The Floquet multiplier, its associated eigenvector, and the spurious eigenvalues from multistep methods are then computed using the periodic QR algorithm [11] with full precision. For reference, we also apply the backward Euler (BE) discretization, which does not produce spurious eigenvalues, to serve as a baseline for convergence comparison.
Through this experiment, we set $\alpha=0.1,\beta=0.1$. The system is discretized over one period into $2^{k}$ points with stepsizes alternating in the ratio $1:2:1:2:$... We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize.
@@
(a) Spurious Eigenvalues of Gear2
(b) Spurious Eigenvalues of Gear3
Figure 2: Convergence Behavior of Spurious Eigenvalues
Figures˜2(a) and 2(b) illustrate the convergence of the spurious eigenvalues associated with Gear2 (two spurious eigenvalues) and Gear3 (four spurious eigenvalues), respectively, as the discretization Stepsize decreases. Since these spurious eigenvalues are extremely small, their logarithms are plotted on the vertical axis, and different solid markers are used to distinguish individual eigenvalues. In addition, based on the multistep method and eq.˜28, we compute the theoretical reference $\nu^{p}$, which represents the expected magnitude of these spurious eigenvalues from the corresponding auxiliary STM, and plot its logarithm as red lines in Figures˜2(a) and 2(b). Strictly speaking, each spurious eigenvalue is equal to the product of the Floquet multiplier of the Spurious system and $\nu^{p}$, where the Floquet multiplier is of moderate size. Therefore, we only compare the order of magnitude, rather than the exact value. Consequently, the slope of the red line corresponds to $\log(\nu)$ and serves as a reference for expected convergence.
The figures show that the computed spurious eigenvalues match the expected magnitude very well, far exceeding the prediction of theorem˜7, which guarantees only linear decay with $\Delta t$.
@@ 7.2 Scenario 1: Coupled Stuart-Landau Oscillators
We fix $\beta=0.5$ and use $\delta$ as the continuation parameter, approaching $\delta=0.25$ from below. Our focus is on the Floquet multipliers of the variational system associated to a limit cycle in eq.˜51, where two oscillators are $\pi$ radians out of phase, i.e., $x_{1}=-x_{2},y_{1}=-y_{2}$. As $\delta$ approaches $0.25$, this limit cycle approaches a heteroclinic loop, and the period grows raplidly, and the Floquet multipliers vary sharply. Our goal is to compute its Floquet multipliers with high accuracy to determine their position relative to the unit circle in the complex plane.
For a given pair of $(\delta,\beta)$, the periodic solution is first computed using existing continuation software—MATCONT [10] in MATLAB and AUTO-07p [26, 9] in FORTRAN. These tools provide not only the periodic orbit but also the Jacobian along the trajectory, enabling us to study the associated variational system.
We test three discretization and eigensolver combinations on the coupled Stuart-Landau system eq.˜51, as listed in table˜2.
@@
Figure˜4 shows the error in computing the oscillatory Floquet multiplier, which should be 1. Both AUTO-07p and the modified MATCONT maintain consistently high accuracy due to their use of fourth-order collocation. In comparison, our Gear4 method performs worse. The reason is as the period increases and the smallest multiplier approaches zero, the spectral gap between Floquet multipliers and spurious eigenvalues decreases. According to theorem˜1, this leads to the separation between Floquet subspace and spurious subspaces decreases, and thus reducing accuracy of each Floquet multipliers. Although the accuracy of the oscillatory multiplier computed by our method is slightly inferior, we are still able to maintain its value close to 1. Since the primary objective is to track the evolution of all four multipliers relative to the unit circle as a parameter varies, the overall performance remains satisfactory, as it still captures all four multipliers reasonably and correctly determines stability.
@@ 7.3 Scenario 2: RF Circuits
In radio frequency (RF) circuits, accurately predicting the phase noise of oscillators is essential for evaluating the performance and stability of oscillatory systems [5]. A critical step in this prediction involves computing the perturbation projection vector (PPV) [5], which corresponds to the eigenvector associated with the dominant Floquet multiplier of an LPTV system. The LPTV system takes the form:
$C(t)^{\mathrm{T}}\frac{\mathrm{d}y}{\mathrm{d}t}+G(t)^{\mathrm{T}}y=0,$ (52)
@@
Here, $C_{11}(t)$ and $G(t)$ are assumed to be nonsingular. The matrix $C_{11}(t)$ is referred to as the capacitance matrix, while $G(t)$ is the conductance matrix. The vector $y_{1}$ represents the part of the nodal voltages, and $y_{2}$ denotes the rest nodes. According to the DAE theory presented in [27], equation eq.˜52 is classified as a semi-explicit index-1 differential-algebraic equation (DAE). It can be systematically decoupled into an ODE and an algebraic constraint:
$\left\{\begin{aligned} &-\frac{\mathrm{d}}{\mathrm{d}t}y_{1}(t)+C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)\right]y_{1}=0,\\ &y_{2}(t)=-G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)y_{1}(t).\end{aligned}\right$. (53)
Our task is to solve the Floquet multiplier closest to 1 and its eigenvectors of the ODE part of eq.˜53, and thus recover the PPV. The physical nature of RF circuits leads to the following three characteristics of eq.˜53:
@@
According to Theorem theorem˜1, the subspace separation between dominant $k$-dimension subspace (the invariant subspace of the dominant $k$ eigenvalues) and the remainder is bounded above by the spectral gap between the $k$-th and $(k+1)$-th largest eigenvalues. Consequently, the dominant one-dimensional subspace is expected to be well isolated, while this separation rapidly decay as more multipliers are included. This expectation is confirmed in fig.˜5(b), indicating this case is a well conditioned case for solving PPV.
@@ 8 Conclusion
A multistep discretization of an LPTV system gives rise to a periodic-PEP for computing Floquet multipliers and their invariant subspaces. This periodic-PEP exhibits clear spectral separation: as the stepsize tends to zero, the dominant eigenvalues converge to the true Floquet multipliers, while the remaining ones decay to zero. By subspace perturbation theory, the computed multipliers and invariant subspaces converge to their exact counterparts at a rate matching the consistency order of the multistep method. Building on this, we propose the pTOAR algorithm to solve all or part of the Floquet eigenpairs efficiently. Numerical experiments demonstrate that the algorithm effectively addresses the two common scenarios, as outlined in the introduction.
Declarations
  • Funding: The research leading to these results received funding from National Key R&D Program of China under Grant Agreement No 2020YFA0711900, 2020YFA0711902.
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use): The authors have no conflicts of interest to declare that are relevant to the content of this article.
  • Ethics approval and consent to participate: Not applicable
  • Consent for publication:Not applicable.
  • Data availability: The datasets generated during and/or analyzed during the current study are available from the corresponding author on reasonable request.
  • Materials availability: Not applicable.
  • Code availability:The code used for the numerical experiments is available from the corresponding author upon reasonable request.
  • Author contribution: Yehao Zhang led the research and contributed to all aspects of this work, including the theoretical development, algorithm design, and numerical experiments. Yuncheng Xu and Yichen Tan participated in the theoretical discussions and provided suggestions for the design of the numerical experiments. Yangfeng Su supervised the research and provided guidance on the overall structure and direction of the study. All authors reviewed and approved the final manuscript.
If any of the sections are not relevant to your manuscript, please include the heading and write ‘Not applicable’ for that section.

### v2

@@ 5 Algorithm and Complexity Analysis
We present pTOAR, an iterative method for computing the dominant $k$ eigenpairs of the pPEP Eq.˜14. It combines the periodic Arnoldi process [18] with TOAR style basis compression [19] to reduce memory and computational cost, making it suitable for large-scale problems. The remainder of this section details the algorithmic steps and its underlying principles.
The periodic Krylov-Schur (pKS) algorithm [18] is well-suited for the linearized pPEP Eq.˜19. It avoids explicitly forming the full product $F_{i+p,i}=L^{(p)}\cdots\allowbreak L^{(1)}$, while maintaining accuracy even for eigenvalues spanning several orders of magnitude. The pKS algorithm constructs a periodic Arnoldi process:
(30) $L^{(i)}V^{(i-1)}=V^{(i)}T^{(i)},$
@@
We explicitly substitute the block form of $L^{(i)}$ into Eq.˜30, and partition the $V^{(i)}$ following the block structure of $L^{(i)}$, as shown in Eq.˜31:
(31) $\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}\begin{bmatrix}V_{1}^{(i-1)}\\ \vdots\\ V_{d-1}^{(i-1)}\\ V_{d}^{(i-1)}\end{bmatrix}=\begin{bmatrix}V_{1}^{(i)}\\ \vdots\\ V_{d-1}^{(i)}\\ V_{d}^{(i)}\end{bmatrix}T^{(i)}.$
The key observation of pTOAR from Eq.˜31 is an information overlap relationship across adjacent basis matrices:
@@
where $v^{(i-1)}$ is the last column of $V^{(i-1)}$, and $v_{j}^{(i-1)}$ is its $j$-th block of $v^{(i-1)}$, also the last column of $V_{j}^{(i-1)}$ in Eq.˜31.
@@
At last, we evaluate memory requirement for storing the input data. For each $L^{(i)}$, it suffices to store the sampled matrices $G^{(i)}$ instead of the coefficient matrix $A_{j}^{(i)}$ in Eq.˜12 and the full $nd\times nd$ matrix $L^{(i)}$. This leads to a memory cost of $n^{\gamma}$ per slice, where $\gamma=1$ if the matrices $A_{j}^{(i)}$ are sparse, and $\gamma=2$ if dense. Combining this with the basis memory and the Ritz extraction memory, the overall memory cost is thus $pn^{\gamma}+pnkd$ under assumption $k\ll n$.
@@
0:Time samples $\{t_{i}\}_{i=0}^{p}$ with $0=t_{0}<t_{1}<\cdots<t_{p}=T$, where $T$ is the period;Discrete evaluations $\{G(t_{i})\}_{i=0}^{p}$ of the system matrix; Target number of eigenpairs $k$.
0:Dominant $k$ Floquet multipliers and their corresponding eigenvectors.
1:Discretization: At each time $t_{i}$, define $\left\{A_{j}^{(i)}\right\}_{j=0}^{d}$ in Eq.˜12 implicitly via matrix-vector products routines, and invoke them to define matrix-vector products of $L^{(i)}$, as inEq.˜18.
2:Eigenpair Computation: Apply the pTOAR to compute the dominant $k$ eigenpairs of the linearized periodic EVP Eq.˜19.
@@ 6 Numerical Results
This section presents three numerical experiments. The first is an artificial example constructed from known Floquet eigenpairs to verify convergence theory and the distribution of parasitic eigenvalues.
The second and third experiments address two common scenarios where Floquet multipliers are needed. In the first scenario, the matrix $G(t;\delta)$ is given explicitly, and tracking the evolution of multipliers with respect to $\delta$ helps detect bifurcations or instability onset [20]. We test this scenario using a benchmark problem and compare our pTOAR method with existing packages MATCONT and AUTO-07p. In the second scenario, $G(t)$ is only available at discrete time samples, without an analytic form; here the focus is often on computing the Floquet eigenvector associated with the dominant multiplier (e.g., in circuit oscillator design) [9, 10]. We illustrate this scenario with an industrial circuit application, where fixed sampling makes collocation infeasible and highlights the advantage of our multistep-based approach.
All numerical experiments are conducted using MATLAB 2024b on a system equipped with an Intel Core i5-1135G7 CPU at 2.4 GHz and 16 GB of RAM.
@@
(32) $e_{\text{val}}=\frac{\left|\lambda-e^{\mu T}\right|}{\left|e^{\mu T}\right|},\quad e_{\text{vec}}=\frac{\|x_{p}-x(T)\|_{2}}{\|x(T)\|_{2}},$
@@
(33) $e_{\text{subspace}}=\Theta(U,V),$
@@
6.1 Convergence Behavior of Floquet multipliers and Parasitic Eiegnvalues
Our first experiment is based on a two-dimensional LTP system given by Eq.˜34,
(34) $\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\alpha\cos^{2}t+(\beta-2)\sin^{2}t&1+(\beta-\alpha-2)\sin t\cos t\\ -1+(\beta-\alpha-2)\sin t\cos t&\alpha\sin^{2}t+(\beta-2)\cos^{2}t\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$.
@@ 6.1 Convergence Behavior of Floquet multipliers and Parasitic Eiegnvalues
(35) $U(t)=\begin{bmatrix}\cos t&\sin t\\ -\sin t&\cos t\end{bmatrix}$.
We discretize one period $0,2\pi]$ with variable stepsize, and apply Gear2, Gear3 methods to obtain the periodic-EVP Eq.˜19. The Floquet multiplier, its associated eigenvector, and the parasitic eigenvalues from multistep methods are then computed using the periodic QR algorithm [5] with full precision. For reference, we also apply the backward Euler (BE) discretization, which does not produce spurio us eigenvalues, to serve as a baseline for convergence comparison.
Through this experiment, we set $\alpha=0.1,\beta=0.1$. The system is discretized over one period into $2^{k}$ points with stepsizes alternating in the ratio $1:2:1:2:$... We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize. Through this experiment, we set $\alpha=0.1$ and $\beta=0.1$. The system is discretized over one period into $2^{k}$ grid points with variable stepsizes. To ensure that the resulting time grid satisfies Assumption (A4), we evenly insert four steps where the stepsize is doubled and four steps where the stepsize is halved, so that the total deviation of the step ratios from unity remains bounded over one period. We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize.
@@
(a) Parasitic Eigenvalues of Gear2
(b) Parasitic Eigenvalues of Gear3
Figure 2: Convergence Behavior of Parasitic Eigenvalues
Figures˜2(a) and 2(b) illustrate the convergence of the parasitic eigenvalues of Gear2 (two eigenvalues) and Gear3 (four eigenvalues), respectively, as the stepsize decreases. Since these eigenvalues are very small, their logarithms are shown, with different markers indicating individual modes. The red lines represent $\log(\nu)$ and serve as a reference for the expected decay rate. As each parasitic eigenvalue equals a Floquet multiplier of moderate magnitude times $\nu^{p}$, we compare only orders of magnitude rather than exact values.
The figures show that the computed parasitic eigenvalues match the expected magnitude very well, far exceeding the prediction of Theorem˜4.1, which guarantees only linear decay with $h$.
@@ 6.2 Scenario 1: Coupled Stuart-Landau Oscillators
We fix $\beta=0.5$ and use $\delta$ as the continuation parameter, approaching $\delta=0.25$ from below. Our focus is on the Floquet multipliers of the variational system associated to a limit cycle in Eq.˜36, where two oscillators are $\pi$ radians out of phase, i.e., $x_{1}=-x_{2},y_{1}=-y_{2}$. As $\delta$ approaches $0.25$, this limit cycle approaches a heteroclinic loop, and the period grows raplidly, and the Floquet multipliers vary sharply. Our goal is to compute its Floquet multipliers with high accuracy to determine their position relative to the unit circle in the complex plane.
For a given pair of $(\delta,\beta)$, the periodic solution is first computed using existing continuation software—MATCONT [11] in MATLAB and AUTO-07p [12] in FORTRAN. These tools provide not only the periodic orbit but also the Jacobian along the trajectory, enabling us to study the associated variational system.
We test three discretization and eigensolver combinations on the coupled Stuart-Landau system Eq.˜36, as listed in Table˜2.
@@
Figure˜4 shows the error in computing the oscillatory Floquet multiplier, which should be 1. Both AUTO-07p and the modified MATCONT maintain consistently high accuracy due to their use of fourth-order collocation. In comparison, our Gear4 method performs worse. The reason is as the period increases and the smallest multiplier approaches zero, the spectral gap between Floquet multipliers and parasitic eigenvalues decreases. This leads to the separation between Floquet subspace and parasitic subspaces decreases, and thus reducing accuracy of each Floquet multipliers. Although the accuracy of the oscillatory multiplier computed by our method is slightly inferior, we are still able to maintain its value close to 1. Since the primary objective is to track the evolution of all four multipliers relative to the unit circle as a parameter varies, the overall performance remains satisfactory, as it still captures all four multipliers reasonably and correctly determines stability.
@@ 6.3 Scenario 2: RF Circuits
In radio frequency (RF) circuits, accurately predicting the phase noise of oscillators is essential for evaluating the performance and stability of oscillatory systems [9]. A critical step in this prediction involves computing the perturbation projection vector (PPV) [9], which corresponds to the eigenvector associated with the dominant Floquet multiplier of an LTP system. The LTP system takes the form:
(37) $C(t)^{\mathrm{T}}\frac{\mathrm{d}y}{\mathrm{d}t}+G(t)^{\mathrm{T}}y=0,$
@@
Here, $C_{11}(t)$ and $G(t)$ are assumed to be nonsingular. The matrix $C_{11}(t)$ is referred to as the capacitance matrix, while $G(t)$ is the conductance matrix. The vector $y_{1}$ represents the part of the nodal voltages, and $y_{2}$ denotes the rest nodes. According to the DAE theory presented in [21], equation Eq.˜37 is classified as a semi-explicit index-1 differential-algebraic equation (DAE). It can be systematically decoupled into an ODE and an algebraic constraint:
(38) $\left\{\begin{aligned} &-\frac{\mathrm{d}}{\mathrm{d}t}y_{1}(t)+C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)\right]y_{1}=0,\\ &y_{2}(t)=-G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)y_{1}(t).\end{aligned}\right$.
Our task is to solve the Floquet multiplier closest to 1 and its eigenvectors of the ODE part of Eq.˜38, and thus recover the PPV. The physical nature of RF circuits leads to the following three characteristics of Eq.˜38:
@@
The subspace separation between dominant $k$-dimension subspace (the invariant subspace of the dominant $k$ eigenvalues) and the remainder is bounded above by the spectral gap between the $k$-th and $(k+1)$-th largest eigenvalues. Consequently, the dominant one-dimensional subspace is expected to be well isolated, while this separation rapidly decay as more multipliers are included. This expectation is confirmed in Fig.˜5(b), indicating this case is a well conditioned case for solving PPV.
@@ 7 Conclusion
A multistep discretization of an LPTV system gives rise to a pPEP for computing Floquet multipliers and their invariant subspaces. This pPEP exhibits clear spectral separation: as the stepsize tends to zero, the dominant eigenvalues converge to the true Floquet multipliers, while the remaining ones decay to zero. By subspace perturbation theory, the computed multipliers and invariant subspaces converge to their exact counterparts at a rate matching the consistency order of the multistep method. Building on this, we propose the pTOAR algorithm to solve all or part of the Floquet eigenpairs efficiently. Numerical experiments demonstrate that the algorithm effectively addresses the two common scenarios, as outlined in the introduction.
Acknowledgments
We would like to acknowledge the assistance of volunteers in putting together this example manuscript and supplement.